OpenStudy (anonymous):

What is the solution to the proportion? 9/x = 4/12

OpenStudy (anonymous):

x = 27, cross multiply. 9 *12 and 4*x. 108=4x, divide 4 on both sides and you get x=27. 9/27 = 1/3 and 4/12 = 1/3

OpenStudy (anonymous):

A. 64% B. 36% C. 26% D. 74%

OpenStudy (anonymous):

I don't understand this one.....

OpenStudy (anonymous):

A. 3 B. 27 C. 13 D. 17

OpenStudy (tkhunny):

Never "cross multiply". It is utterly silly and useless. Simply solve for the variable. 9/x = 4/12 Simplfy where possible 9/x = 1/3 x cannot be zero, so multiply by that. 9 = x/3 Multiply by 3. x = 27 Done!